A triangle has a measure of its side X+3, X+8, 2X+3. If the perimeter of the same triangle is 46 units, what is the measure of its shortest side?

Answer by nerdybill(7384)

A triangle has a measure of its side X+3, X+8, 2X+3. If the perimeter of the same triangle is 46 units, what is the measure of its shortest side?
.
Shortest side would be x+3
.
we need to find out what x is... apply what we know of the perimeter:
X+3 + X+8 + 2X+3 = 46
4X+3 + 8 + 3 = 46
4X+14 = 46
4X = 32
x = 8
.
shortest side:
x+3 = 8+3 = 11 units